OpenStudy (anonymous):

a stone thrown from ground level return to same level 4s after. with what speed was the stone thrown? take g=10ms-2?

OpenStudy (anonymous):

please solve it for me

OpenStudy (anonymous):

Use s=ut+at^2/2 Put s=0 , a=-10 , t=4

OpenStudy (anonymous):

This is a simple kinematics problem. Separate the stone's flight into two sections. One going up, one going down. They both take the same amount of time, meaning two seconds up, two down. The velocity at the top of the stone's flight is zero! Then, use d = Vot + 1/2 a*t^2, starting from the second half of the stone's flight, from when it stops moving at the top to the bottom of its flight path. With the new distance, plug into Vf^2 = Vo^2 + 2ad. That should give you the starting velocity. (Vf is the "starting" velocity because we already assigned Vo to zero, the speed at the top of the flight path.)